3. Let a: R+R be the regular curve defined by a(t) = (213, 121, 61). Find B, K and .
Answer #1

2: R3 R3 24 )= (2 t? 12t, 6+²) Tit)= x 2 (4) unit tangent vector Ñita I (t) Unit normal vects 1 7 ll F (*) xÑ( do BilarialT(4) = (4 42). 24 - £.(2 toon + (4²42) 0 - 2 12 t) } (t + 2)2 (+²+2) 2 (422). 2-24 ( 24 tuh + ² + 2)2 6 + ² +4 +- 2 ts) i +42 43 î tî hat + 2+2) th (-22²-4t) B (t) = (2x + 4 214) X = 6+ 12 126 - î [144-0] } [72¢² - 144x2] th Co-1444] = nuuê +72+2kit)= 3 (+²+2) 2 ora or kate |T164) = 2/470 b 6(+²+2) 3(+²42)3 112 ll TE 2 (t). (2 (+) x 2 (A)) 12 (4) x 2 (A)/2 d . 12